# Correspondence-Free, Function-Based Sim-to-Real Learning for Deformable Surface Control

### v1

@@ III Differentiable Alignment for Training
To enable correspondence-free training of the function-based sim-to-real network, we introduce a confidence-map weighted Chamfer distance as a robust metric for evaluating discrepancies between the predicted shape $\mathcal{S}_{j}^{*}$ and the captured shape $\mathcal{S}_{j}^{p}$. This weighted Chamfer distance serves as the primary loss term for aligning $\mathcal{S}_{j}^{*}$ with $\mathcal{S}_{j}^{p}$ during training. The process jointly optimizes the function-prediction network $\mathcal{N}_{rbf}$ and the confidence-prediction network $\mathcal{N}_{conf}$, aided by additional regularization terms – namely, confidence regularization, normal compatibility, and geometric regularization. Note that we assume the coordinate systems of $\mathcal{S}_{j}^{*}$ and $\mathcal{S}_{j}^{p}$ have been pre-aligned through rigid registration before taking the sim-to-real training. Finally, we describe how this sim-to-real learning framework can be adapted to scenarios involving partially captured markers with known correspondences.
@@ III-A Confidence Weighted Chamfer Distance
Chamfer distance has been widely used in previous works for rigid registration (e.g., [27]). In contrast, our task involves non-rigid registration on models with large deformations, which presents additional challenges. Notably, the captured shapes of free-form surfaces undergoing large deformations often contain missing regions. To address this, we introduce a real-valued confidence map $w_{c}(\cdot)\in[0,1]$ on the surface of each simulated shape $\mathcal{S}_{j}$, representing the likelihood of finding reliable correspondences on the scanned shape $\mathcal{S}_{j}^{p}$.
Without loss of generality, this confidence map is parameterized by a neural network $\mathcal{N}_{conf}$, whose parameters are optimized during the sim-to-real training process. The input to $\mathcal{N}_{conf}$ includes a surface point $\mathbf{p}\in\mathcal{S}$ and the set of control points $\mathcal{S}^{c}(\mathbf{a})$ that encode the simulated shape, allowing the network to adapt to varying shape configurations. The network outputs a scalar in the range $[0,1]$, representing the confidence that the point $\mathbf{p}$ has a reliable correspondence on the scanned shape $\mathcal{S}_{j}^{p}$ – i.e., $w_{c}(\mathbf{p})=\mathcal{N}_{conf}(\mathbf{p},\mathcal{S}^{c}(\mathbf{a}))$.
@@
$+\frac{1}{|\mathcal{S}_{j}^{p}|}\sum_{\mathbf{x}\in\mathcal{S}^{p}_{j}}\min_{\mathbf{p}^{*}\in\mathcal{S}_{j}^{*}}\|\mathbf{p}^{*}-\mathbf{x}\|_{2}^{2},$ (3)
where $\mathbf{p}^{*}$ is the sim-to-real transferred surface point corresponding to $\mathbf{p}=\mathbf{B}(u_{p},v_{p},\mathcal{S}^{c}(\mathbf{a}))$, as defined in Eq. (2). Here, $|\cdot|$ denotes the number of sample points on surface.
@@ III-B Loss Functions for Training
The loss function for shape alignment based training is defined according to the confidence weighted Chamfer distance as
@@
Besides of this primary loss, other three loss terms are needed to serve for purpose of regularization.
Figure 5: When there is a large gap between (a) real and (b) simulated shapes, the RBF based space warping may generate a weird shape as shown in (c). This problem can be effectively solved by introducing the function compatibility requirement as a geometric regularization loss during training – see (d) for the result of sim-to-real correction with this addtional regularization. The color map illustrates the distribution of geometric errors.
@@ III-B2 Normal Compatibility
To ensure that the predicted surface $\mathcal{S}_{j}^{*}$ has normals compatible with those of the captured surface $\mathcal{S}_{j}^{p}$, we introduce a normal compatibility loss:
@@ III-B3 Geometric Regularization
$\mathcal{L}_{gr}$ $:=\sum_{j}\|\sum_{i=1}^{N}\bm{\beta}_{i}\|^{2}+\sum_{j}\|\sum_{i=1}^{N}\bm{\beta}_{i}(\mathbf{q}_{i})_{x}\|^{2}$ (8)
@@
The effectiveness of this geometric regularization term has been demonstrated in Fig.5(d).
@@
which combines the aforementioned loss terms with corresponding weights. The function-prediction network $\mathcal{N}_{rbf}$ and the confidence-prediction network $\mathcal{N}_{conf}$ are jointly optimized by minimizing $\mathcal{L}_{total}$ (see also Fig.2 for the illustration of training pipeline). The weight coefficients $\omega_{cr}=12.0$, $\omega_{nc}=0.5$, and $\omega_{gr}=\text{1.5e-4}$ are empirically selected based on experiments and remain fixed across all examples.
@@ III-C Training with Known Correspondences
  • Replace the nearest-neighbor search in Eq.(III-A) with the known correspondence term $\|\mathbf{p}_{i}^{*}-\mathbf{x}_{i}\|_{2}^{2}$;
@@
The effectiveness of this simplified sim-to-real training process using a MoCap system will be demonstrated in Sec.V.
@@ IV Neural Networks Based Kinematic Computing
This section presents the neural network-based computational pipeline for kinematics. Shape control of deformable free-form surfaces can be realized by our fast IK solution.
Figure 6: NN-based pipeline for forward kinematic computing.
@@ IV-A NN-based Forward Kinematics
While the recently developed fast simulator provides an efficient solution to predict the deformed shape of soft robots [29], applying it to estimate gradients of the IK objective by numerical differentiation remains time-consuming [18]. Here we propose to employ an NN-based computational pipeline for forward kinematics. The input of our forward kinematic network is the actuation parameter $\mathbf{a}$, and the output is the control points $\{\mathcal{S}^{c}\}$ as a compact representation of freeform surfaces.
Given a set of randomly generated actuation parameters as $\{\mathbf{a}_{k}\}$, we can run the numerical simulators to obtain the deformed shapes $\{\mathcal{S}\}$ and generate their corresponding control points as $\{\mathcal{S}^{c}_{k}\}$. A training dataset with $M$ such pairs of results $\{\mathbf{a}_{k}:\mathcal{S}^{c}_{k}\}_{k=1,\ldots,M}$ can be obtained by the simulator using a reduced physical model (e.g., [29]). It is important to generate sample shapes that span the entire space of shape variation. The dataset is then employed to train a network $\mathcal{N}_{fk}$ that can give the fast simulation result with an input actuation as $\mathcal{S}^{c}=\mathcal{N}_{fk}(\mathbf{a})$.
The gap between the simplified simulation and the physical reality can be well captured and corrected by our sim-to-real network. As a result, for any point $(u_{p},v_{p})$ sampled in the $u,v$-parametric domain, its position on the physical model can be predicted by Eq.(2). This NN-based computational pipeline for forward kinematics was illustrated in Fig.6. The predicted shape $\mathcal{S}^{*}(\mathbf{a})$ can be obtained as a set of points $\{\mathbf{p}^{*}\}$, which is a differentiable function in terms of the actuation.
@@ IV-B Gradient Iteration Based Inverse Kinematics
The IK computation is formulated as an optimization problem that minimizes the shape difference between the predicted shape $\mathcal{S}^{*}$ and the target shape $\mathcal{S}^{t}$. That is
$\arg\min_{\mathbf{a},\mathbf{R},\mathbf{t}}D(\mathcal{S}^{*}(\mathbf{a}),\mathcal{S}^{t})=\sum_{\mathbf{p}^{*}\in\mathcal{S}^{*}(\mathbf{a})}\|\mathbf{p}^{*}-\left(\mathbf{R}\mathbf{c}^{*}+\mathbf{t}\right)\|^{2}$ (11)
where $\mathbf{c}^{*}$ is the closet point of $\mathbf{p}^{*}$ on the target shape $\mathcal{S}^{t}$. The rotation matrix $\mathbf{R}$ and the translation vector $\mathbf{t}$ are applied to the target model to eliminate the influence of orientation variation.
The objective function can be minimized using the gradient descent method with linear search as illustrated in Fig.7, where the pseudo-code of our IK solver can be found in Appendix A. Thanks to the formulation of NN-based forward kinematics and our function based sim-to-real learning, the gradients of $D(\cdot)$ including sim-to-real transfer can be computed analytically. That is
@@
$\mathbf{Q}$ denotes those terms contributed by the set of kernels for RBF-base warping while $\mathbf{B}$ is for the surface point $\mathbf{p}$. Detail formulas can be found in Appendix B.
After each iteration of updating $\mathbf{a}$, we apply an ICP-based rigid registration [30] to determine a new pair of $(\mathbf{R},\mathbf{t})$. In other words, the optimization process alternates between updating the actuation parameter $\mathbf{a}$ and adjusting the pose of target model $(\mathbf{R},\mathbf{t})$. However, in certain scenarios – such as the soft manipulator examples presented in this paper – this registration step becomes unnecessary, as the pose of the target shape remains fixed. In such cases, we simply set $(\mathbf{R}=\mathbf{I},\mathbf{t}=\mathbf{0})$ in the relevant equations.
@@ V Implementation Details and Results
We have implemented the proposed approach in C++ and Python. The network training phase is implemented on the PyTouch platform with the learning rate at $0.001$ and the maximum number of epochs as $150$. For the inference phase, we transferred the trained networks to integrate with our IK solver in C++ running on CPU. The analytical gradients of the networks are employed in our IK solver. Our source code will be released on GitHub upon the acceptance of this paper. All the training and computational tests are conducted on a PC with Intel i7-12700H CPU, RTX $3060$ GPU and $32$ GB RAM.
@@ V-A4 Compressible Soft Manipulator
The four hardware setups were selected to span representative deformation modes and structural complexities in pneumatically actuated soft robots. The deformable membrane provides inflation-induced deformations with regular chamber geometry, while the robotic mannequin involves inflation through freeform chamber boundaries with overlaps, resulting in more complex and spatially coupled surface deformations. The two soft manipulators highlight structural bending, with one driven purely by bending and the other combining inflation / compression and bending, thereby introducing hybrid deformation behaviors. Together, these systems cover the spectrum from simple to complex chamber inflation and from pure bending to coupled bending–inflation, demonstrating that our correspondence-free sim-to-real learning framework generalizes across diverse actuation principles and deformation characteristics rather than being tailored to a single morphology.
@@
 Kernels are selected coincident with the positions of 34 real markers for MoCap input on the robotic mannequin; but kernels are sampled in a decoupled way in the $u,v$-domain for other examples.
@@ V-B2 Function prediction network $\mathcal{N}_{rbf}$
For learning the sim-to-real transfer, we employ a network architecture with 2 hidden layers as $\mathcal{N}_{rbf}$, where each layer has a certain number of neurons using the ReLU activation function. The input layer contains the positions of $30\times 30$ control points, and the output layer is the coefficients for the RBF-based warping function as $\bm{\gamma}\in\mathbb{R}^{3(N+4)}$ with $N$ being the number of kernels. The training of $\mathcal{N}_{rbf}$ is taken on shapes captured by either 3D scanner or MoCap while randomly varying the pneumatic actuation within the working range of the chambers. Among these captured shapes, a certain number of frames have missing regions (or markers). However, all shapes are employed to train $\mathcal{N}_{rbf}$. The total number of valid samples (or markers) for different cases have been reported in Table II. Using the dataset of deformable mannequin generated from MoCap as an eaxmple, it can be observed from the study as shown in Fig. 9(b) that the prediction error starts to increase when the total number of neurons exceeds 50 – i.e., overfitting happens. Based on this analysis, we select 24 neurons for each layer in $\mathcal{N}_{rbf}$.

### v2

@@ III Differentiable Alignment for Training
To enable correspondence-free training of the function-prediction network $\mathcal{N}_{rbf}$, we introduce a confidence-map weighted Chamfer distance as a robust metric for evaluating discrepancies between the predicted shape $\mathcal{S}_{j}^{*}$ and the captured shape $\mathcal{S}_{j}^{p}$. This weighted Chamfer distance serves as the primary loss for aligning $\mathcal{S}_{j}^{*}$ with $\mathcal{S}_{j}^{p}$ during training. The process jointly optimizes $\mathcal{N}_{rbf}$ and a newly introduced confidence-prediction network $\mathcal{N}_{conf}$, aided by additional regularization terms – namely, confidence regularization, normal compatibility, and geometric regularization. Finally, we describe how this framework can be adapted to scenarios involving partially captured markers with known correspondences.
@@ III-A Confidence Weighted Chamfer Distance
Chamfer distance has been widely used in previous works for rigid registration (e.g., [27]). In contrast, our task involves non-rigid registration on models with large deformations, which presents additional challenges. Moreover, the captured shapes of free-form surfaces undergoing large deformations often contain missing regions. To address this, we introduce a real-valued confidence map $w_{c}(\cdot)\in[0,1]$ on the surface of each simulated shape $\mathcal{S}_{j}$, representing the likelihood of finding reliable correspondences on the scanned shape $\mathcal{S}_{j}^{p}$.
Without loss of generality, this confidence map is parameterized by a neural network $\mathcal{N}_{conf}$, whose parameters are determined jointly during the sim-to-real training process. The input to $\mathcal{N}_{conf}$ includes a surface point $\mathbf{p}\in\mathcal{S}$ and the set of control points $\mathcal{S}^{c}(\mathbf{a})$ that encode the simulated shape, allowing the network to adapt to varying shape configurations. The network outputs a scalar in the range $[0,1]$, representing the confidence that the point $\mathbf{p}$ has a reliable correspondence on the scanned shape $\mathcal{S}_{j}^{p}$ – i.e., $w_{c}(\mathbf{p})=\mathcal{N}_{conf}(\mathbf{p},\mathcal{S}^{c}(\mathbf{a}))$.
@@
$+\frac{1}{|\mathcal{S}_{j}^{p}|}\sum_{\mathbf{x}\in\mathcal{S}^{p}_{j}}\min_{\mathbf{p}^{*}\in\mathcal{S}_{j}^{*}}\|\mathbf{x}-\mathbf{p}^{*}\|_{2}^{2},$ (6)
where $\mathbf{p}^{*}\in\mathbb{R}^{3}$ is the sim-to-real transferred surface point corresponding to $\mathbf{p}=\mathbf{B}(u_{p},v_{p},\mathcal{S}^{c}(\mathbf{a}))$, as defined in Eq. (4). Here, $|\cdot|$ denotes the number of sample points on surface. As only the scanned shape $\mathcal{S}^{p}_{j}$ may have missed regions, the confidence map is not applied to the closest point search from the physically sampled geometry $\mathcal{S}^{p}_{j}$ to the predicted shape $\mathcal{S}^{*}_{j}$ that is assumed to be complete.
@@ III-B Loss Functions for Training
The loss function for shape alignment based training is defined by the confidence weighted Chamfer distances as
@@
Besides of this primary loss, other three loss terms are defined to serve for purpose of regularization.
Figure 5: When there is a large gap between (a) real and (b) simulated shapes, the RBF based space warping may generate a weird shape as shown in (c). This problem can be effectively solved by introducing the function compatibility requirement as a geometric regularization during training – see (d) for the sim-to-real result with this additional regularization. The color map illustrates the distribution of geometric errors.
@@ III-B2 Normal Compatibility
To ensure the normals of a predicted surface $\mathcal{S}_{j}^{*}$ compatible with those of the captured surface $\mathcal{S}_{j}^{p}$, we introduce a normal compatibility loss:
@@ III-B3 Geometric Regularization
$\mathcal{L}_{gr}$ $:=\sum_{j}\|\sum_{i=1}^{N}\bm{\beta}_{i}\|^{2}+\sum_{j}\|\sum_{i=1}^{N}\bm{\beta}_{i}(\mathbf{q}_{i})_{x}\|^{2}$
@@
The effectiveness of this geometric regularization term has been demonstrated in Fig.5(c, d).
@@
which combines the aforementioned loss terms with corresponding weights. The function-prediction network $\mathcal{N}_{rbf}$ and the confidence-prediction network $\mathcal{N}_{conf}$ are jointly trained by minimizing $\mathcal{L}_{total}$ (see also Fig.2 for the illustration of training pipeline). The weight coefficients $\omega_{cr}=12.0$, $\omega_{nc}=0.5$, and $\omega_{gr}=\text{1.5e-4}$ are empirically selected based on experiments and remain fixed across all examples.
@@ III-C Training with Known Correspondences
  • Replace the nearest-neighbor search in Eq.(III-A) with the known correspondence term $\|\mathbf{p}_{i}^{*}-\mathbf{x}_{i}\|_{2}^{2}$, which gives
    $\tilde{D}(\mathcal{S}_{j}^{*},\mathcal{S}_{j}^{p})=\frac{1}{|\mathcal{S}_{j}|}\sum_{\mathbf{p}_{i}\in\mathcal{S}_{j}}w_{c}(\mathbf{p}_{i})\|\mathbf{p}_{i}^{*}-\mathbf{x}_{i}\|_{2}^{2};$ (13)
@@
The effectiveness of this simplified sim-to-real training process using a MoCap system will be demonstrated in Sec.V (see the right of Fig.13(b)).
@@ IV Neural Networks Based Kinematic Computing
This section presents the NN-based computational pipeline for kinematics. With the help of a sim-to-real network trained above, accurate shape control of deformable free-form surfaces can be realized by our fast IK solution.
Figure 6: NN-based pipeline for forward kinematic computing, by which the accurate shape of a deformable robot represented as a set of 3D point $\{\mathbf{p}^{*}\in\mathbb{R}^{3}\}$ can be predicted by the input actuation parameters as a vector $\mathbf{a}$.
Input: The actuation vector $\mathbf{a}$.
1
Output: The deformed shape $\mathcal{S}^{*}$ as a set of points $\{\mathbf{p}^{*}\}$.
2
3 Compute the simulated shape as control pnts. $\mathcal{S}^{c}$ by Eq.(14);
4 Determine the coefficients of RBF function as $\bm{\gamma}$ by Eq.(3);
5for every sample point $(u_{p},v_{p})$ do
6 Compute its position as $\mathbf{p}=\mathbf{B}(u_{p},v_{p},\mathcal{S}^{c})$ by Eq.(1);
7 Update its position as $\mathbf{p}^{*}=\Phi_{\bm{\gamma}}(\mathbf{p})$ by Eq.(4);
8 end for
9
10return $\mathcal{S}^{*}=\{\mathbf{p}^{*}\}$;
ALGORITHM 1 NN-Based Forward Kinematics
@@ IV-A NN-based Forward Kinematics
While the recently developed fast simulator provides an efficient solution to predict the deformed shape of soft robots [30], applying it to estimate gradients of the IK computation by numerical differentiation remains time-consuming [18]. To enable fast IK computation, we propose to employ an NN-based computational pipeline for forward kinematics as shown in Fig.6. The input of our shape prediction network $\mathcal{N}_{fk}$ is the actuation parameter $\mathbf{a}$, and the output is the control points $\{\mathcal{S}^{c}\}$ as a compact representation of freeform surfaces. This gives
$\mathcal{S}^{c}=\mathcal{N}_{fk}(\mathbf{a}).$ (14)
The network $\mathcal{N}_{fk}$ actually serves as a surrogate model for the numerical simulation of soft robot’s deformation.
The network can be trained via a supervised learning process. Given a set of randomly generated actuation parameters as $\{\mathbf{a}_{k}\}$, we can run the numerical simulators to obtain the deformed shapes and generate their corresponding control points as $\{\mathcal{S}^{c}_{k}\}$. A training dataset with $M$ such pairs of results $\{\mathbf{a}_{k}:\mathcal{S}^{c}_{k}\}_{k=1,\ldots,M}$ can be obtained by a simulator (e.g., [30]). It is important to generate example shapes that span the entire space of shape variation. The dataset is then employed to train the network $\mathcal{N}_{fk}$.
The computation of forward kinematics is conducted with the help of $\mathcal{N}_{fk}$ and $\mathcal{N}_{s2r}$. Given an actuation $\mathbf{a}$, the control points of its simulated surface $\mathcal{S}^{c}$ can be obtained by Eq.(14). The corresponding coefficients $\bm{\gamma}$ of the RBF function for sim-to-real transfer can then be obtained by Eq.(3). As a result, for any point $(u_{p},v_{p})$ sampled in the $u,v$-parametric domain, its position on the physical model can be predicted by Eq.(4). This NN-based computational pipeline for forward kinematics was illustrated in Fig.6 – see also the pseudo-code given in Algorithm 1. The predicted shape $\mathcal{S}^{*}(\mathbf{a})$ is represented by a set of points $\{\mathbf{p}^{*}\}$, which is a differentiable function in terms of the actuation $\mathbf{a}$.
1
Output: The optimized actuation $\mathbf{a}_{opt}$.
2
3Set $i=0$ and $r_{0}=1.0$;
6while $i<i_{max}$ and $r_{i}>\lambda$ do
 /* Update transformation matrix */
 /* Gradient-based iteration with fixed transformation matrix */
8   Evaluate the objective function $D_{i}=D(\mathbf{a}_{i})$;
9  Compute the gradient of $D(\mathbf{a}_{i})$ as $\frac{\partial D}{\partial\mathbf{a}}$ by Eq.(16);
10  Compute the optimal step size $h$ by line search [31];
11  Set $\mathbf{a}_{i+1}=\mathbf{a}_{i}-h\frac{\partial D}{\partial\mathbf{a}}$ and $i=i+1$;
12  Evaluate the objective function $D_{i}=D(\mathbf{a}_{i})$;
[…]
14 end while
15return $\mathbf{a}_{opt}=\mathbf{a}_{i}$;
ALGORITHM 2 Fast Inverse Kinematics Solver
@@ IV-B Gradient Iteration Based Inverse Kinematics
Following the strategy proposed in [18], the IK computation is formulated as an optimization problem that minimizes the shape difference between the predicted shape $\mathcal{S}^{*}$ and the target shape $\mathcal{S}^{t}$ while allowing the pose change. That gives
$\arg\min_{\mathbf{a},\mathbf{R},\mathbf{t}}D(\mathcal{S}^{*}(\mathbf{a}),\mathcal{S}^{t})=\sum_{\mathbf{p}^{*}\in\mathcal{S}^{*}(\mathbf{a})}\|\mathbf{p}^{*}-\left(\mathbf{R}\mathbf{c}^{*}+\mathbf{t}\right)\|^{2},$ (15)
where $\mathbf{c}^{*}$ is the closet point of $\mathbf{p}^{*}$ on the target shape $\mathcal{S}^{t}$. The rotation matrix $\mathbf{R}$ and the translation vector $\mathbf{t}$ are applied to the target model to eliminate the influence of pose change. The objective function can be minimized using the gradient descent method with linear search. After each iteration of updating $\mathbf{a}$, we apply an ICP-based rigid registration [32] to determine a new pair of $(\mathbf{R},\mathbf{t})$. In other words, the optimization process alternates between updating the actuation parameter $\mathbf{a}$ and adjusting the pose of target model $(\mathbf{R},\mathbf{t})$ [^2] For cases where the target shape’s pose is fixed (e.g., the soft manipulator examples), the registration step is omitted by setting $(\mathbf{R}=\mathbf{I},\mathbf{t}=\mathbf{0})$.. The diagram of our fast IK solver is as illustrated in Fig.7, and the pseudo-code of our IK solver can be found in Algorithm 2.
Thanks to the formulation of NN-based forward kinematics including sim-to-real, the gradients of $D(\cdot)$ in Eq.(15) can be computed analytically. That is
@@
$\mathbf{Q}$ denotes those terms contributed by the set of kernels for RBF-base warping while $\mathbf{B}$ is for the surface point $\mathbf{p}$. Detail formulas can be found in Appendix A-A.
@@ V Implementation Details and Results
We have implemented the proposed approach in C++ and Python. The network training phase is implemented on the PyTouch platform with the learning rate at $0.001$ and the maximum number of epochs as $150$. For the inference phase, we transferred the trained networks to integrate with our IK solver in C++ running on CPU. The analytical gradients of the networks are employed in our IK solver. Our source code will be released on GitHub upon the acceptance of this paper at: https://yinggwan.github.io/CFS2R.github.io. All the training and computational tests are conducted on a PC with Intel i7-12700H CPU, RTX $3060$ GPU and $32$ GB RAM.
@@ V-A4 Compressible Soft Manipulator
The four hardware setups were selected to span representative deformation modes and structural complexities in pneumatically actuated soft robots. The deformable membrane provides inflation-induced deformations with regular chamber geometry, while the robotic mannequin involves inflation through freeform chamber boundaries with overlaps, resulting in more complex and spatially coupled surface deformations. The two soft manipulators highlight structural bending, with one driven purely by bending and the other combining inflation / compression and bending, thereby introducing hybrid deformation behaviors. Together, these systems cover the spectrum from simple to complex chamber inflation and from pure bending to coupled bending, inflation and compression, demonstrating that our correspondence-free sim-to-real learning framework generalizes across diverse actuation principles and deformation characteristics rather than being tailored to a single morphology.
@@
 When Mocap is employed for the soft mannequin example, the kernels are located to be consistent with 34 real markers; for other examples, kernels are sampled in a decoupled way in the $u,v$-domain.
@@ V-B2 Function prediction network $\mathcal{N}_{rbf}$
For learning the sim-to-real transfer, we employ a network architecture with 2 or 3 hidden layers as $\mathcal{N}_{rbf}$, where each layer has a certain number of neurons using the ReLU activation function. The input layer contains the positions of $30\times 30$ control points, and the output layer is the coefficients for the RBF-based warping function as $\bm{\gamma}\in\mathbb{R}^{3(N+4)}$ with $N$ being the number of kernels. The training of $\mathcal{N}_{rbf}$ is taken on shapes captured by either 3D scanner or MoCap while randomly varying the pneumatic actuation within the working range of the chambers. Among these captured shapes, a certain number of frames have missing regions (or markers). However, all shapes are employed to train $\mathcal{N}_{rbf}$. The total number of valid samples (or markers) for different cases have been reported in Table II. Using the dataset of deformable mannequin generated from MoCap as an eaxmple, it can be observed from the study as shown in Fig. 9(b) that the prediction error starts to increase when the total number of neurons exceeds 50 – i.e., overfitting happens. Based on this analysis, we select 24 neurons for each layer in $\mathcal{N}_{rbf}$.